# TimeLens: Rethinking Video Temporal Grounding with Multimodal LLMs

### v1

@@ 1 Introduction
Recent multimodal large language models (MLLMs) have excelled at understanding “what” happens in a video, yet they largely fail when asked “when.” This limitation is central to the task of video temporal grounding (VTG). The challenge is twofold: 1) VTG necessitates a fundamental shift from coarse semantic aggregation to fine-grained time-aware perception; 2) Distinguishing queried events requires modeling long-term visual dynamics over appearance-centric features, which are notoriously difficult to annotate and learn. As MLLMs become integral to perception [yuan2021closerlookat, wu2025numberit, perception_test, vcr_bench] and reasoning systems [nagrani2025minerva, nagrani2024neptune, liu2025videomind, chain_of_frames, zhang2025vitcot, longvila_rl], equipping them with robust temporal awareness is no longer optional, but essential [vtg_survey, univtg, timechat, wang2025timer1, team2025vidi].
@@
This work focuses on post-training MLLMs with leading temporal grounding ability. This investigation is a straightforward extension given the recent progress in pretrained foundation MLLMs [qwen2-5-vl, Qwen3-VL, internvideo2]. Different from heavily studied general understanding tasks, recipes for fine-grained grounding tasks are not yet to be established. This paper aims to systematically investigate core components of building time-aware MLLMs (Fig. 1) along two primary dimensions: data quality and algorithmic design.
Our investigation starts by exposing critical flaws in evaluation benchmarks. We find that existing VTG benchmarks [charades-sta, activitynet-captions, qvhighlights] not only lack a clear comparison between leading proprietary and open-source models but are also rife with low-quality queries and erroneous timestamps. This noisy data may render current leaderboards misleading and misguide research efforts. To rectify this, we undertook a meticulous data overhaul. We first defined strict criteria for query and timestamp quality, in terms of uniqueness, existence, clarity, and accuracy. We then manually re-annotated three popular datasets (Charades-STA [charades-sta], ActivityNet Captions [activitynet-captions], QVHighlights [qvhighlights]) to create TimeLens-Bench, a rigorously cross-validated benchmark. As shown in Fig. 2(a), the necessity of this correction is confirmed by a dramatic re-ranking of models on TimeLens-Bench compared to their performance on legacy benchmarks, proving the unreliability of prior evaluation standards. Beyond evaluation, we also fix the noisy training data by automated re-annotation, yielding TimeLens-100K, a large-scale, high-quality training dataset.
With our curated data suite as a solid foundation, we conduct in-depth explorations on the algorithmic design principles from three key aspects. First, for timestamp representation, we discovered that a simple yet effective interleaved textual encoding strategy outperforms more complex alternatives. Second, we determined that VTG is fundamentally a perception-driven task, and thus employ a pure thinking-free reinforcement learning with verifiable rewards (RLVR) approach that outperforms other training paradigms in both efficiency and performance. Finally, our detailed analysis of RLVR training reveals two key recipes for both performance and training efficiency: (1) early stopping when reward metrics plateau, and (2) difficulty-based data sampling. By integrating these insights and design principles, we ultimately develop TimeLens models, a family of MLLMs with superior VTG capability. As shown in Fig. 2(b), our model achieves state-of-the-art performance among open-source models and even surpasses proprietary models such as GPT-5 and Gemini-2.5-Flash.
@@ 2 Related Work
Numerous VTG datasets have been proposed, spanning diverse domains [qvhighlights, activitynet-captions, oncescu2021queryd, wang2024cosmo_cap, huang2024vtimellm-internvid-vtime, grauman2022ego4d, tacos]. Early works [charades-sta, 2d-tan, lu2019debug] trained and evaluated models on the training and test splits of a single benchmark [tacos, activitynet-captions] to assess their ability to fit single-domain data distribution. In recent works [liu2025videomind, timechat, guo2024trace], large diverse corpuses composed of multiple different source datasets  [didemo, oncescu2021queryd, wang2024cosmo_cap, huang2024vtimellm-internvid-vtime, zala2023hirest, liu2024etbench] are aggregated for training, and a suite of distinct benchmarks [activitynet-captions, charades-sta, qvhighlights] are used to probe the models’ real-world cross-domain generalizability.
@@
Substantial works focus on algorithmic designs to improve MLLMs’ VTG capability. One line of research explores model architectures, including token compression methods to reduce computation on long videos  [timechat, zeng2024timesuite] and timestamp encoding strategies to align the timestamps of each frame with its corresponding features [chen2024timemarker, wu2025numberit, li2025unitime, ge2025arc-hunyuan-video, zeng2025distime]. Another line of works investigate training strategies: introducing various supervised fine-tuning tasks to improve VTG performance [zeng2024timesuite, cheng2025tempura], or designing verifiable rewards to improve performance via reinforcement learning [wang2025timer1, tvg-r1, li2025tempsamp_r1, yue2025tempo_r0].
@@ 3.2 Manual Auditing and Refinement
Upon completion of each small data batch, every sample is assigned to a different annotator for cross-validation and error correction. If the error rate in a batch exceeds a threshold, the entire batch is rejected for re-annotation and then validated again. For annotator selection and training, we sampled a small subset of data for trial annotation with over a dozen vendors, then selected the vendor with the highest quality and consistency. Before formal annotation, we provided a detailed handbook and conducted several training sessions. The annotation interface and detailed manual are provided in Appendix E of the appendix.
@@
GPT-4o [hurst2024gpt-4o] 60.6 44.5 23.5 41.8 55.2 41.4 25.8 40.4 69.0 54.8 38.5 52.1
GPT-5 [OpenAI2025_GPT5] 59.3 42.0 22.0 40.5 57.4 44.9 30.4 42.9 72.4 60.4 46.4 56.8
Gemini-2.0-Flash [comanici2025gemini-2-5] 66.4 53.5 27.1 46.7 62.9 54.0 37.7 49.3 76.2 66.4 48.3 60.8
Gemini-2.5-Flash [comanici2025gemini-2-5] 68.7 56.1 30.6 48.6 66.8 57.5 41.3 52.5 78.2 69.4 55.0 64.3
Gemini-2.5-Pro [comanici2025gemini-2-5] 74.1 61.1 34.0 52.8 72.3 64.2 47.1 58.1 84.1 75.9 61.1 70.4
@@
VideoChat-Flash-7B [li2024videochat_flash] 60.2 37.9 17.8 39.7 35.5 21.8 10.5 24.8 45.2 30.6 16.7 32.7
VideoChat-R1-7B [li2025videochat_r1] 51.9 30.8 11.7 33.7 35.0 23.9 11.3 25.0 29.3 19.1 9.4 21.5
Time-R1-7B [wang2025timer1] 57.9 32.0 16.9 36.6 44.8 31.0 19.0 33.1 65.8 51.5 36.1 49.2
MiMo-VL-7B [mimo-vl] 57.9 42.6 20.5 39.6 49.3 38.7 22.4 35.5 57.1 42.6 28.4 41.5
@@ 3.3 Empirical Analysis on TimeLens-Bench
In this section, we present our efforts and findings by applying the above annotation pipeline to existing datasets. We focus on three most widely-used temporal grounding benchmarks: Charades-STA [charades-sta], ActivityNet Captions [activitynet-captions], and QVHighlights [qvhighlights]. These datasets exhibit diversity across video domains, video durations, and query semantics. They are all manually annotated and generally considered the highest-quality VTG datasets available. Therefore, analyzing them offers a representative view of the quality and issues prevalent in existing data. Through diagnosis and refinement, we release TimeLens-Bench, comprising refined versions of the three aforementioned benchmarks: Charades-TimeLens, ActivityNet-TimeLens, and QVHighlights-TimeLens. Together, they form a comprehensive evaluation suite that combines diversity with high quality. Detailed statistics for these benchmarks are provided in Appendix A.
@@
We evaluate various frontier models on both the original and refined benchmarks, observing drastically contrasting performance trends. As illustrated in Fig. 2(a), on the original benchmarks, we observe a surprising phenomenon: frontier proprietary models like Gemini-2.5-Pro [comanici2025gemini-2-5] receive poor scores, whereas open-source models [wang2025timer1, qwen2-5-vl] attain significantly higher ones. Conversely, on our refined benchmarks, this trend reverses. The proprietary models exhibit much better results, though with room for improvement, while the open-source models suffer a substantial performance degradation, lagging far behind their proprietary counterparts. This reversal indicates that the original benchmarks produce misleading results due to inherent quality flaws, while our refined benchmarks yield results that align more closely with real-world user experience, providing reliable evaluation for developing better VTG models.
@@ 3.4 Training Data Re-annotation
By applying our manual pipeline from Sec. 3.2 to a sampled subset of existing VTG training corpus [didemo, oncescu2021queryd, wang2024cosmo_cap, huang2024vtimellm-internvid-vtime, zala2023hirest], we found that the training data exhibits an even higher error rate compared to the evaluation benchmarks. This motivated us to refine training data based on scalable re-annotation. Given the vast scale of the training sets, we employ an automated pipeline to improve their quality based on advanced multimodal models. Owing to the poor quality of these training datasets, especially the high proportion of queries that fail to meet our criteria in Sec. 3.1, we re-annotate the videos rather than refining existing labels. Through this process, we curate TimeLens-100K, a large-scale, high-quality, and diverse VTG training set. Additional details are provided in Appendix F.
@@ 4 Benchmarking Grounding MLLMs
As shown in Tab. 1, we observe a significant performance gap between existing open-source and proprietary models, and our TimeLens models substantially narrow this gap. TimeLens-7B delivers substantial improvements over its baseline, demonstrating the effectiveness of the insights and best practices obtained from our experiments in Sec. 5. It surpasses strong open-source competitors such as Time-R1-7B [wang2025timer1] and MiMo-VL-7B [mimo-vl], as well as proprietary models like GPT-4o [hurst2024gpt-4o] and GPT-5 [OpenAI2025_GPT5]. More remarkably, on the already stronger baseline Qwen3-VL-8B, our TimeLens-8B model still achieves substantial performance gains, establishing a new state-of-the-art among open-source models and even surpassing frontier proprietary models like Gemini-2.5-Flash [comanici2025gemini-2-5].
@@ 5 Exploring Algorithmic Designs
Experimental Setup. Our experiments use Qwen2.5-VL-7B [qwen2-5-vl] as the baseline. For RLVR experiments, we employ GRPO [shao2024deepseekmath] as optimization method. We use TimeLens-Bench for evaluation and TimeLens-100K for training. To ensure rigor, all ablation studies are based on the final, best-performing model configuration, isolating the impact of a single design choice at a time. Due to limited computational resources, we adopt a lower per-frame resolution for our ablation experiments. More implementation details are provided in Appendix B of the appendix.
@@ 5.1 Timestamp Encoding
  • Position-embedding based. These methods adapt position embeddings in LLMs to represent the temporal position of each frame. For example, MRoPE [qwen2-5-vl, mimo-vl] and 3D RoPE [team2025kwai_keye] extend pure-text RoPE to multimodal scenarios, encoding the spatial and temporal dimensions of video frame tokens.
  • Visual overlay. These methods [ge2025arc-hunyuan-video, wu2025numberit, cheng2025tempura] directly overlay timestamps or frame index onto each frame, enabling MLLMs to “read” the temporal position through their OCR capabilities.
  • Textual encoding. These methods convert timestamps into text tokens using the MLLM’s text tokenizer. There are two main variants: the Interleaved approach [guo2025seed1_5_vl, li2025unitime, yao2025genS, hong2024cogvlm2, chen2024timemarker] in Fig. 5(a) inserts timestamp tokens before the visual tokens of each frame. In contrast, the Non-interleaved approach [li2024videochat_flash, wang2025spacevllm, li2023videochat] adds an instruction like “This video samples $N$ frames of a $T$-second video at $t_{1},t_{2},\ldots$ seconds.” into the prompt.
@@ 5.2 Optimization Paradigms
Earlier works [timechat, huang2024vtimellm-internvid-vtime, zeng2024timesuite, guo2024trace, guo2025vtg-llm] employ supervised fine-tuning (SFT) to improve MLLMs’ VTG capability. Recently, some works [wang2025timer1, tvg-r1] utilize reinforcement learning with verifiable rewards (RLVR), following a “think-then-answer” approach [guo2025deepseek_r1] (details in Appendix B): during sampling, the model first generates an explicit thinking process and then produces the final answer. The task-specific VTG accuracy reward is computed only on the final answer. Despite these efforts, there lacks a systematic comparison of the respective merits of these methods, leaving some key questions unanswered:
  • Is RLVR superior to SFT? While the pioneering work Time-R1 [wang2025timer1] demonstrates that RLVR outperforms SFT, they compare the two methods using the same amount of training data, despite RLVR requiring significantly more training time. A fair comparison under equal training budgets remains absent.
  • Is explicit “thinking” necessary for RLVR? Recent works suggest that the thinking process is not essential when applying RLVR to visual perception such as counting [li2025cls-rl, perception_test]. Whether this holds for VTG, a predominantly perception-oriented task, remains unanswered.
  • Does a preceding SFT phase benefit RLVR? An SFT phase prior to RLVR is typically employed to enhance the model’s capability and facilitate subsequent RLVR training [team2025kwai_keye, mimo-vl]. However, whether this preceding SFT phase actually improves final performance in the VTG scenario remains unexplored.
@@ 5.3 Recipes for RLVR Training
How long should we train? In SFT, the prevailing wisdom is “train longer, generalize better” [hoffer2017train_longer]. Given training data with sufficient scale and quality, we typically train MLLMs for at least one full epoch over the entire dataset, ensuring the model sees as much data as possible to enhance generalization. However, whether this strategy is optimal for RLVR remains to be explored.
@@
How to sample training data? For RLVR training, it is crucial to select samples with appropriate difficulty relative to the model, and many works propose assessing training sample difficulty and employing difficulty-aware sampling [yuan2025vl_cogito, wang2025timer1, hong2025glm_4_5]. To evaluate the impact of sample difficulty on video temporal grounding, we conduct experiments on our TimeLens-100K high-quality training corpus. Following prior works [wang2025timer1, yuan2025vl_cogito], we use the model to be trained to perform offline inference on the training data, compute IoU metrics to estimate sample difficulty, and then perform Gaussian sampling based on sample difficulty (details in Appendix B). By varying the mean of the Gaussian distribution, we obtain training sets with different difficulty levels relative to the model, and conduct RLVR training on each set independently.
@@
6 Conclusion

### v2

@@ 1 Introduction
Recent multimodal large language models (MLLMs) have excelled at understanding “what” happens in a video, yet they largely fail when asked “when.” This limitation is central to the task of video temporal grounding (VTG). The challenge is twofold: 1) VTG necessitates a fundamental shift from coarse semantic aggregation to fine-grained time-aware perception; 2) Distinguishing queried events requires modeling long-term visual dynamics over appearance-centric features, which are notoriously difficult to annotate and learn. As MLLMs become integral to perception [58, 55, 43, 44] and reasoning systems [39, 40, 36, 13, 66, 6], equipping them with robust temporal awareness is no longer optional, but essential [26, 35, 46, 53, 49].
@@
This work focuses on post-training MLLMs with leading temporal grounding ability. This investigation is a straightforward extension given the recent progress in pretrained foundation MLLMs [3, 2, 54]. Different from heavily studied general understanding tasks, recipes for fine-grained grounding tasks are not yet to be established. This paper aims to systematically investigate core components of building time-aware MLLMs (Fig. 1) along two primary dimensions: data quality and algorithmic design.
Our investigation starts by exposing critical flaws in evaluation benchmarks. We find that existing VTG benchmarks [11, 25, 27] not only lack a clear comparison between leading proprietary and open-source models but are also rife with low-quality queries and erroneous timestamps. This noisy data may render current leaderboards misleading and misguide research efforts. To rectify this, we undertook a meticulous data overhaul. We first defined strict criteria for query and timestamp quality, in terms of uniqueness, existence, clarity, and accuracy. We then manually re-annotated three popular datasets (Charades-STA [11], ActivityNet Captions [25], QVHighlights [27]) to create TimeLens-Bench, a rigorously cross-validated benchmark. As shown in Fig. 2(a), the necessity of this correction is confirmed by a dramatic re-ranking of models on TimeLens-Bench compared to their performance on legacy benchmarks, proving the unreliability of prior evaluation standards. Beyond evaluation, we also fix the noisy training data by automated re-annotation, yielding TimeLens-100K, a large-scale, high-quality training dataset.
With our curated data suite as a solid foundation, we conduct in-depth explorations on the algorithmic design principles from three key aspects. First, for timestamp representation, we discover that a simple yet effective interleaved textual encoding strategy outperforms more complex alternatives. Second, we determine that VTG is fundamentally a perception-driven task, and thus employ a pure thinking-free reinforcement learning with verifiable rewards (RLVR) approach that outperforms other training paradigms in both efficiency and performance. Finally, our detailed analysis of RLVR training reveals two key recipes for both performance and training efficiency: (1) early stopping when reward metrics plateau, and (2) difficulty-based data sampling. By integrating these insights and design principles, we ultimately develop TimeLens models, a family of MLLMs with superior VTG capability. As shown in Fig. 2(b), our model achieves state-of-the-art performance among open-source models and even surpasses proprietary models such as GPT-5 and Gemini-2.5-Flash.
@@ 2 Related Work
Numerous VTG datasets have been proposed, spanning diverse domains [27, 25, 41, 50, 22, 14, 45]. Early works [11, 65, 38] trained and evaluated models on the training and test splits of a single benchmark [45, 25] to assess their ability to fit single-domain data distribution. In recent works [36, 46, 18], large diverse corpuses composed of multiple different source datasets  [1, 41, 50, 22, 60, 37] are aggregated for training, and a suite of distinct benchmarks [25, 11, 27] are used to probe the models’ real-world cross-domain generalizability.
@@
Substantial works focus on algorithmic designs to improve MLLMs’ VTG capability. One line of research explores model architectures, including token compression methods to reduce computation on long videos  [46, 61], timestamp encoding strategies to align the timestamps of each frame with its corresponding features [5, 55, 34, 12, 62, 51, 28]. Another line of works investigate training strategies: introducing various supervised fine-tuning tasks to improve VTG performance [61, 7], or designing verifiable rewards to improve performance via reinforcement learning [53, 4, 33, 59].
@@ 3.2 Manual Auditing and Refinement
Upon completion of each small data batch, every sample is assigned to a different annotator for cross-validation and error correction. If the error rate in a batch exceeds a threshold, the entire batch is rejected for re-annotation and then validated again. For annotator selection and training, we sampled a small subset of data for trial annotation with over a dozen vendors, then selected the vendor with the highest quality and consistency. Before formal annotation, we provided a detailed handbook and conducted several training sessions. The annotation interface and detailed manual are provided in Appendix G of the appendix.
@@
GPT-4o [23] 60.6 44.5 23.5 41.8 55.2 41.4 25.8 40.4 69.0 54.8 38.5 52.1
GPT-5 [42] 59.3 42.0 22.0 40.5 57.4 44.9 30.4 42.9 72.4 60.4 46.4 56.8
Gemini-2.0-Flash [8] 66.4 53.5 27.1 46.7 62.9 54.0 37.7 49.3 76.2 66.4 48.3 60.8
Gemini-2.5-Flash [8] 68.7 56.1 30.6 48.6 66.8 57.5 41.3 52.5 78.2 69.4 55.0 64.3
Gemini-2.5-Pro [8] 74.1 61.1 34.0 52.8 72.3 64.2 47.1 58.1 84.1 75.9 61.1 70.4
@@
VideoChat-Flash-7B [31] 60.2 37.9 17.8 39.7 35.5 21.8 10.5 24.8 45.2 30.6 16.7 32.7
VideoChat-R1-7B [32] 51.9 30.8 11.7 33.7 35.0 23.9 11.3 25.0 29.3 19.1 9.4 21.5
Time-R1-7B [53] 57.9 32.0 16.9 36.6 44.8 31.0 19.0 33.1 65.8 51.5 36.1 49.2
TRACE [18] 37.2 21.8 9.6 27.1 43.4 33.9 22.0 32.7 49.7 39.1 28.1 39.0
TRACE-uni [18] 38.2 22.9 10.4 28.1 44.3 35.1 22.6 33.6 49.9 40.0 29.2 39.8
TimeSuite [61] 56.3 35.5 18.0 38.1 27.1 17.5 8.6 19.8 27.1 16.9 9.9 21.7
Grounded-VideoLLM [51] 43.3 28.7 13.5 30.0 39.2 29.6 19.5 30.0 43.7 33.8 22.5 33.4
MiMo-VL-7B [9] 57.9 42.6 20.5 39.6 49.3 38.7 22.4 35.5 57.1 42.6 28.4 41.5
@@ 3.3 Empirical Analysis on TimeLens-Bench
In this section, we present our efforts and findings by applying the above annotation pipeline to existing datasets. We focus on three most widely-used temporal grounding benchmarks: Charades-STA [11], ActivityNet Captions [25], and QVHighlights [27]. These datasets exhibit diversity across video domains, video durations, and query semantics. They are all manually annotated and generally considered the highest-quality VTG datasets available. Therefore, analyzing them offers a representative view of the quality and issues prevalent in existing data. Through diagnosis and refinement, we release TimeLens-Bench, comprising refined versions of the three aforementioned benchmarks: Charades-TimeLens, ActivityNet-TimeLens, and QVHighlights-TimeLens. Together, they form a comprehensive evaluation suite that combines diversity with high quality. Detailed statistics for these benchmarks are provided in Appendix B.
@@
We evaluate various frontier models on both the original and refined benchmarks, observing drastically contrasting performance trends. As illustrated in Fig. 2(a), on the original benchmarks, we observe a surprising phenomenon: frontier proprietary models like Gemini-2.5-Pro [8] receive poor scores, whereas open-source models [53, 3] attain significantly higher ones. Conversely, on our refined benchmarks, this trend reverses. The proprietary models exhibit much better results, though with room for improvement, while the open-source models suffer a substantial performance degradation, lagging far behind their proprietary counterparts. This reversal indicates that the original benchmarks produce misleading results due to inherent quality flaws, while our refined benchmarks yield results that align more closely with real-world user experience, providing reliable evaluation for developing better VTG models.
@@ 3.4 Training Data Re-annotation
By applying our manual pipeline from Sec. 3.2 to a sampled subset of existing VTG training corpus [1, 41, 50, 22, 60], we found that the training data exhibits an even higher error rate compared to the evaluation benchmarks. This motivated us to refine training data based on scalable re-annotation. Given the vast scale of the training sets, we employ an automated pipeline to improve their quality based on advanced multimodal models. Owing to the poor quality of these training datasets, especially the high proportion of queries that fail to meet our criteria in Sec. 3.1, we re-annotate the videos rather than refining existing labels. Through this process, we curate TimeLens-100K, a large-scale, high-quality, and diverse VTG training set. Additional details are provided in Appendix H.
@@ 4 Benchmarking Grounding MLLMs
As shown in Tab. 1, we observe a significant performance gap between existing open-source and proprietary models, and our TimeLens models substantially narrow this gap. TimeLens-7B delivers substantial improvements over its baseline, demonstrating the effectiveness of the insights and best practices obtained from our experiments in Sec. 5. It surpasses strong open-source competitors such as Time-R1-7B [53] and MiMo-VL-7B [9], as well as proprietary models like GPT-4o [23] and GPT-5 [42]. More remarkably, on the already stronger baseline Qwen3-VL-8B, our TimeLens-8B model still achieves substantial performance gains, establishing a new state-of-the-art among open-source models and even surpassing frontier proprietary models like Gemini-2.5-Flash [8].
@@ 5 Exploring Algorithmic Designs
Experimental Setup. Our experiments use Qwen2.5-VL-7B [3] as the baseline. For RLVR experiments, we employ GRPO [47] as optimization method. We use TimeLens-Bench for evaluation and TimeLens-100K for training. To ensure rigor, all ablation studies are based on the final, best-performing model configuration, isolating the impact of a single design choice at a time. Due to limited computational resources, we adopt a lower per-frame resolution for our ablation experiments. More implementation details are provided in Appendix C of the appendix.
@@ 5.1 Timestamp Encoding
  • Position-embedding based. These methods adapt position embeddings in LLMs to represent the temporal position of each frame. For example, MRoPE [3, 9] and 3D RoPE [48] extend pure-text RoPE to multimodal scenarios, encoding the spatial and temporal dimensions of video frame tokens.
  • Visual overlay. These methods [12, 55, 7] directly overlay timestamps or frame index onto each frame, enabling MLLMs to “read” the temporal position through their OCR capabilities.
  • Textual encoding. These methods convert timestamps into text tokens using the MLLM’s text tokenizer. There are two main variants: the Interleaved approach [16, 34, 56, 20, 5] in Fig. 5(a) inserts timestamp tokens before the visual tokens of each frame. In contrast, the Non-interleaved approach [31, 52, 29] adds an instruction like “This video samples $N$ frames of a $T$-second video at $t_{1},t_{2},\ldots$ seconds.” into the prompt.
@@ 5.2 Optimization Paradigms
Earlier works [46, 22, 61, 18, 17] employ supervised fine-tuning (SFT) to improve MLLMs’ VTG capability. Recently, some works [53, 4] utilize reinforcement learning with verifiable rewards (RLVR), following a “think-then-answer” approach [15] (details in Appendix C): during sampling, the model first generates an explicit thinking process and then produces the final answer. The task-specific VTG accuracy reward is computed only on the final answer. Despite these efforts, there lacks a systematic comparison of the respective merits of these methods, leaving some key questions unanswered:
  • Is RLVR superior to SFT? While the pioneering work Time-R1 [53] demonstrates that RLVR outperforms SFT, they compare the two methods using the same amount of training data, despite RLVR requiring significantly more training time. A fair comparison under equal training budgets remains absent.
  • Is explicit “thinking” necessary for RLVR? Recent works suggest that the thinking process is not essential when applying RLVR to visual perception such as counting [30, 43]. Whether this holds for VTG, a predominantly perception-oriented task, remains unanswered.
  • Does a preceding SFT phase benefit RLVR? An SFT phase prior to RLVR is typically employed to enhance the model’s capability and facilitate subsequent RLVR training [48, 9]. However, whether this preceding SFT phase actually improves final performance in the VTG scenario remains unexplored.
@@ 5.3 Recipes for RLVR Training
How long should we train? In SFT, the prevailing wisdom is “train longer, generalize better” [19]. Given training data with sufficient scale and quality, we typically train MLLMs for at least one full epoch over the entire dataset, ensuring the model sees as much data as possible to enhance generalization. However, whether this strategy is optimal for RLVR remains to be explored.
@@
How to sample training data? For RLVR training, it is crucial to select samples with appropriate difficulty relative to the model, and many works propose assessing training sample difficulty and employing difficulty-aware sampling [57, 53, 21]. To evaluate the impact of sample difficulty on video temporal grounding, we conduct experiments on our TimeLens-100K high-quality training corpus. Following prior works [53, 57], we use the model to be trained to perform offline inference on the training data, compute IoU metrics to estimate sample difficulty, and then perform Gaussian sampling based on sample difficulty (details in Appendix C). By varying the mean of the Gaussian distribution, we obtain training sets with different difficulty levels relative to the model, and conduct RLVR training on each set independently.
@@
Acknowledgements. This work is supported by the National Key R&D Program of China (No. 2022ZD0160900), the Basic Research Program of Jiangsu (No. BK20250009), the Fundamental and Interdisciplinary Disciplines Breakthrough Plan of the Ministry of Education of China (No. JYB2025XDXM118), and the Collaborative Innovation Center of Novel Software Technology and Industrialization.
Supplementary Material
Appendix A Conclusion